# EDNet: A Distortion-Agnostic Speech Enhancement Framework with Gating Mamba Mechanism and Phase Shift-Invariant Training

### v1

@@ Abstract
Speech signals in real-world environments are frequently affected by various distortions such as additive noise, reverberation, and bandwidth limitation, which may appear individually or in combination. Traditional speech enhancement methods typically rely on either masking, which focuses on suppressing non-speech components while preserving observable structure, or mapping, which seeks to recover clean speech through direct transformation of the input. Each approach offers strengths in specific scenarios but may be less effective outside its target conditions. We propose the Erase and Draw Network (EDNet), a distortion-agnostic speech enhancement framework designed to handle a broad range of distortion types without prior assumptions about task or input characteristics. EDNet consists of two main components: (1) the Gating Mamba (GM) module, which adaptively combines masking and mapping through a learnable gating mechanism that selects between suppression (Erase) and reconstruction (Draw) based on local signal features, and (2) Phase Shift-Invariant Training (PSIT), a shift tolerant supervision strategy that improves phase estimation by enabling dynamic alignment during training while remaining compatible with standard loss functions. Experimental results on denoising, dereverberation, bandwidth extension, and multi distortion enhancement tasks show that EDNet consistently achieves strong performance across conditions, demonstrating its architectural flexibility and adaptability to diverse task settings.
Index Terms:
Speech enhancement, denoising, dereverberation, bandwidth extension, distortion-agnostic, shift-tolerant phase modeling
@@ I Introduction
In real-world conditions, speech signals are frequently corrupted by diverse acoustic distortions such as additive noise, reverberation, and bandwidth limitations. These degradations often co-occur in unpredictable combinations, severely hindering both perceptual intelligibility and the performance of downstream applications [1, 2, 3]. While speech enhancement (SE) systems essentially aim to restore clean speech from such degraded inputs, most existing approaches are optimized for a specific type of distortion. This task-specific design paradigm may fail to reflect the reality of practical deployment environments, where distortion types are neither known in advance nor mutually exclusive, and may even vary over time.
Conventional SE methods can be largely categorized into masking-based and mapping-based approaches. Masking-based methods suppress non-speech components while preserving speech structure by predicting a mask, making them effective for tasks such as denoising [4, 5, 6, 7, 8, 9, 10, 11] and dereverberation [12, 13, 14, 15, 16]. However, their reliance on input spectral cues limits their ability to reconstruct fully lost speech components [17]. In contrast, mapping-based methods directly transform degraded speech into clean speech, facilitating the reconstruction of missing components—especially in tasks such as bandwidth extension [18, 19, 20, 21, 22, 23]. Despite these strengths, these methods can excessively alter preserved speech content, undermining identity preservation and leading to suboptimal objective scores [24, 25, 23]. These contrasting behaviors highlight the inherent trade-off between preservation and reconstruction in existing SE strategies. Several hybrid approaches [26, 27, 28, 29] have been proposed to leverage the complementary strengths of masking and mapping, aiming to balance selective suppression with generative reconstruction. These methods have shown effectiveness when tailored to specific enhancement tasks, but most integrate the two methods through fixed fusion schemes—such as summation or averaging—that are agnostic to the characteristics of the input. While such strategies can work well in certain cases, applying a static combination regardless of distortion type or task may fall short of optimality and often requires manual design choices about how the fusion should be performed.
@@
Complementing this architectural flexibility on the magnitude side, we also address the orthogonal challenge of phase reconstruction. Accurate phase estimation is crucial for perceptual quality, particularly in conditions involving reverberation or multiple interacting distortions [30]. However, phase learning is inherently difficult due to its periodicity and sensitivity to temporal shifts [31]. Conventional phase reconstruction methods [32, 33, 34, 31] typically constrain training to a single ground truth (GT) phase per sample, implicitly requiring the model to align its output exactly with the GT phase. This strict constraint can be unnecessary, as minor phase shifts are, from a signal processing perspective, nearly indistinguishable from small temporal shifts in the time domain, and therefore have negligible impact on perceived speech quality [35]. While some prior works [36, 35] attempt to address these issues through data augmentation or specialized objectives, such methods often sacrifice compatibility with standard loss functions or require non-standard training pipelines.
@@
  • We propose EDNet, a distortion-agnostic speech enhancement framework designed to handle diverse and compound distortions by embedding structural adaptability into its architecture.
@@ II-A Masking-based SE
Masking-based methods operate primarily in the time-frequency (TF) domain by learning masks that selectively suppress non-speech components while preserving speech-dominant regions. Early works like the Ideal Binary Mask (IBM) [6, 7, 12, 13] apply hard decisions to each TF unit, significantly improving intelligibility but introducing artifacts due to coarse quantization [37, 38]. To address this, Ideal Ratio Mask (IRM) [8, 9, 14] was proposed, enabling soft masks with continuous values between 0 and 1 for smoother enhancement. The Spectral Magnitude Mask (SMM) [39] further relaxes the range constraint, allowing unbounded mask values and increasing modeling flexibility. As phase information gained importance, methods such as Phase-Sensitive Mask (PSM) [1, 40] and Complex Ideal Ratio Mask (cIRM) [41, 15, 16] were developed. Masking has been also extended to the latent feature space in time-domain architectures such as TasNet [42], where masks are applied directly to learned feature representations, avoiding the constraints of the STFT domain. This feature-level masking has become central in many state-of-the-art models, particularly in the context of speech separation [43, 44]. Despite their wide adoption, masking-based models remain fundamentally limited in scenarios involving severely missing spectral information, such as bandwidth limitation and packet loss [17].
@@ II-B Mapping-based SE
Mapping-based speech enhancement methods aim to learn a nonlinear transformation that reconstructs clean speech directly from noisy input, typically in the waveform or spectral domain. Although masking remains the dominant approach in denoising, several studies have explored mapping-based alternatives. A representative example is SEGAN [4], which employs adversarial training for waveform mapping. Subsequent works increased generator capacity [45] and combined adversarial loss with spectral-domain losses [46] for improved stability and fidelity. In dereverberation, mapping-based methods are motivated by the need of modeling long-range temporal smearing introduced by room reverberation. Early spectral regression models [47] used fully connected networks to predict clean spectra, followed by U-Net structures [48], and variants that replaced skip connections with multiple convolutional modules to improve feature propagation [49]. More recently, deformable convolutional networks (DCN) [50] have been introduced context-sensitive transformations that extend beyond conventional masking and exhibit characteristics of both filtering and reconstruction-based approaches. Bandwidth extension is an area where mapping dominates, as it requires inferring missing high-frequency content. TF-domain methods [19, 20, 22] estimate high-band magnitude spectra from narrowband inputs and use heuristics or learned strategies for phase reconstruction. Time-domain approaches [18, 21] directly predict wideband waveforms, while vododer-based method such as NVSR [23] uses a two-stage pipeline of spectral mapping followed by waveform generation via a neural vocoder. Despite their generative flexibility, mapping-based models can inadvertently distort clean input regions due to their global transformation behavior [24, 25, 23].
@@ II-C Hybrid Approachs
To overcome the individual limitations of masking and mapping approaches, several studies have explored hybrid architectures that combine the two, aiming to take advantage of their complementary strengths: masking offers selective noise suppression, while mapping enables detailed signal reconstruction. PhaseDCN [26] applies the hybrid approach with a focus on improving phase modeling to enhance denoising performance. It first generates a masked magnitude spectrum using an IRM and concatenates it with the noisy input. The combined features are fed into a complex-valued mapping network that produces a final output, implementing early stage fusion at the feature level. DBDIUNet [27] focuses on noisy and reverberant conditions by predicting a cIRM and a magnitude spectrogram through dual decoders, then averaging their outputs to produce the final result. The masking-and-inpainting [28] method targets low signal-to-noise ratios (SNR) and non-stationary noise by first removing highly corrupted time-frequency regions using a binary mask, then reconstructing the removed region with a convolutional inpainting module. These two stages are executed sequentially, without an explicit fusion step. CMGAN [29], designed to handle various SE tasks such as denoising, dereverberation, and bandwidth extension, predicts magnitude masks and complex spectrograms in parallel and combines them using element-wise summation. This demonstrates the potential of hybrid methods in building distortion-agnostic enhancement systems.
Despite differences in design and application, these approaches share several limitations. Most rely on fixed fusion strategies—such as averaging, concatenation, or summation—that do not adapt to the specific characteristics of the input. Moreover, these fusion mechanisms are typically global, lacking the ability to modulate the region-wise contributions of masking and mapping dynamically across time and frequency. These observations underscore the need for more flexible integration strategies that can coordinate masking and mapping in a content-aware, task-agnostic manner.
@@ II-D Phase Modeling
Early speech enhancement models primarily focused on magnitude estimation [51, 52], often reusing the noisy phase during synthesis. However, it is now widely recognized that accurate phase reconstruction is essential for achieving high perceptual quality [30]. One line of research addresses phase prediction by directly operating in the complex domain either through masking [53, 54, 55] or mapping [56, 57], where both magnitude and phase are implicitly modeled via real and imaginary components. While conceptually straightforward, this approach often suffers from optimization instability and limited interpretability, especially when paired with magnitude-focused objectives such as spectral magnitude loss or mask estimation targets [10]. To improve the stability and effectiveness of phase modeling, recent studies [11, 10, 22] have moved toward explicitly decoupling magnitude and phase and modeling them through separate branches.
Supervision for phase modeling varies by modeling approach. In complex-valued formulations, losses are typically applied in terms of complex spectral distance. In contrast, explicit phase prediction often requires trigonometric losses [32] or anti-wrapping functions [33] to handle phase discontinuities, and is frequently combined with derivative-based objectives such as group delay (GD) or instantaneous angular frequency (IAF) [34, 31]. Most of these methods, however, assume a single ground-truth (GT) phase, implicitly enforcing strict alignment, which can penalize perceptually irrelevant shifts and destabilize learning [35].
While the limitations of strict phase alignment have begun to attract attention, only a few studies have explored ways to relax this constraint. PhaseAug [36], originally proposed for vocoder training, introduces random phase shifts during training to expose the model to multiple plausible alignments. However, the loss is still computed with respect to a single shifted target, meaning the model is implicitly encouraged to align with the shifted version of the GT. Ku et al. [35] take a different approach by removing phase GT entirely and using magnitude-phase consistency as an indirect supervision signal. While this avoids explicit alignment, it prevents the use of standard loss functions—such as waveform or complex spectral losses—since applying them would implicitly reintroduce supervision toward a fixed target phase. These issues suggest the need for a supervision strategy that resolves alignment ambiguity while preserving compatibility with existing loss functions and training pipelines.
@@ III-A Dual-stream Architecture
EDNet is designed with a dual-stream architecture to separately process magnitude and phase spectrograms, as shown in  Fig. 1. Instead of using a shared encoder, it employs separate dilated DenseNet encoders [10] with feature channel dimensions of 64 and 32 for magnitude and phase, respectively. This separation is intended to allow magnitude and phase, which have distinct characteristics—magnitude follows a linear scale, where phase exhibits periodicity—to be processed individually, aiming to prevent interference between their estimations. The decoder also follows this structure, using separate dilated DenseNet decoders [10] for magnitude and phase. The magnitude decoder directly reconstructs the spectrogram using ReLU activation [58], as feature-wise masking is handled within the GM module.
The input speech signal is first transformed using Short-Time Fourier Transform (STFT), decomposing it into magnitude and phase spectrograms. The magnitude spectrogram undergoes power-law compression to enhance feature learning. Each spectrogram is then encoded separately using the respective encoders. Noisy magnitude features are progressively refined through three stacked GM modules. The enhanced magnitude features are then concatenated with the noisy phase features, reduced to 32 channels via a 1×1 convolution, and processed through three stacked Time-Frequency (TF) Mamba blocks [11], which model long-range dependencies across temporal and spectral domains using bidirectional Mamba [59], to refine the phase. Finally, the enhanced spectrograms are reconstructed using the decoders, and the enhanced waveform is obtained by combining them with Inverse STFT (ISTFT).
@@ III-C1 Phase Shift Estimation
where $f_{AW}$ is an anti-wrapping function that ensures phase differences remain within a range $[-\pi,\pi]$. However, the presence of the anti-wrapping function introduces non-linearity, making it difficult to derive a closed-form solution for the optimal $n$.
@@
This formulation allows us to approximate the optimal shift by assuming a locally linear relationship between the anti-wrapped phase difference and the shift term $2\pi fn/N$. For a restricted range of $n$ in $[-0.5,0.5]$, where shift term remains within the anti-wrapping range, this approximation holds effectively, enabling a differentiable solution. Taking the derivative with respect to $n$ and solving for zero, we obtain:

### v2

# EDNet: A Versatile Speech Enhancement Framework with Gating Mamba Mechanism and Phase Shift-Invariant Training

@@ Abstract
Speech signals in real-world environments are frequently affected by various distortions such as additive noise, reverberation, and bandwidth limitation, which may appear individually or in combination. Traditional speech enhancement methods typically rely on either masking, which focuses on suppressing non-speech components while preserving observable structure, or mapping, which seeks to recover clean speech through direct transformation of the input. Each approach offers strengths in specific scenarios but may be less effective outside its target conditions. We propose the Erase and Draw Network (EDNet), a versatile speech enhancement framework designed to handle a broad range of distortion types without prior assumptions about task or input characteristics. EDNet consists of two main components: (1) the Gating Mamba (GM) module, which adaptively combines masking and mapping through a learnable gating mechanism that selects between suppression (Erase) and reconstruction (Draw) based on local signal features, and (2) Phase Shift-Invariant Training (PSIT), a shift tolerant supervision strategy that improves phase estimation by enabling dynamic alignment during training while remaining compatible with standard loss functions. Experimental results on denoising, dereverberation, bandwidth extension, and multi distortion enhancement tasks show that EDNet consistently achieves strong performance across conditions, demonstrating its architectural flexibility and adaptability to diverse task settings.
@@ I Introduction
In real-world conditions, speech signals are frequently corrupted by diverse acoustic distortions such as additive noise, reverberation, and bandwidth limitations. These degradations often co-occur in unpredictable combinations, severely hindering both perceptual intelligibility and the performance of downstream applications [weninger2015speech, desjardins2014effect, kwak2024voxmm]. While speech enhancement (SE) systems essentially aim to restore clean speech from such degraded inputs, most existing approaches are optimized for a specific type of distortion. This task-specific design paradigm may fail to reflect the reality of practical deployment environments, where distortion types are neither known in advance nor mutually exclusive, and may even vary over time.
Conventional SE methods can be largely categorized into masking-based and mapping-based approaches. Masking-based methods suppress non-speech components while preserving speech structure by predicting a mask, making them effective for tasks such as denoising [pascual2017segan, yin2022tridentse, hu2001speech, yilmaz2004blind, narayanan2013ideal, srinivasan2006binary, lu2023explicit, chao2024investigation] and dereverberation [roman2011intelligibility, may2014generalization, li2017ideal, kothapally2022skipconvgan, williamson2017time]. However, their reliance on input spectral cues limits their ability to reconstruct fully lost speech components [chao2025universal]. In contrast, mapping-based methods directly transform degraded speech into clean speech, facilitating the reconstruction of missing components—especially in tasks such as bandwidth extension [kuleshov2017audio, abel2018simple, li2015deep, wang2021towards, lu2024towards, liu2022neural]. Despite these strengths, these methods can excessively alter preserved speech content, undermining identity preservation and leading to suboptimal objective scores [nossier2020mapping, kumar2020nu, liu2022neural]. These contrasting behaviors highlight the inherent trade-off between preservation and reconstruction in existing SE strategies. Several hybrid approaches [zhang2021deep, zhang2023dual, hao2020masking, abdulatif2024cmgan] have been proposed to leverage the complementary strengths of masking and mapping, aiming to balance selective suppression with generative reconstruction. These methods have shown effectiveness when tailored to specific enhancement tasks, but most integrate the two methods through fixed fusion schemes—such as summation or averaging—that are agnostic to the characteristics of the input. While such strategies can work well in certain cases, applying a static combination regardless of distortion type or task may fall short of optimality and often requires manual design choices about how the fusion should be performed.
@@
Complementing this architectural flexibility on the magnitude side, we also address the orthogonal challenge of phase reconstruction. Accurate phase estimation is crucial for perceptual quality, particularly in conditions involving reverberation or multiple interacting distortions [paliwal2011importance]. However, phase learning is inherently difficult due to its periodicity and sensitivity to temporal shifts [masuyama2020phase]. Conventional phase reconstruction methods [takamichi2018phase, ai2023neural, zhang2024unrestricted, masuyama2020phase] typically constrain training to a single ground truth (GT) phase per sample, implicitly requiring the model to align its output exactly with the GT phase. This strict constraint can be unnecessary, as minor phase shifts are, from a signal processing perspective, nearly indistinguishable from small temporal shifts in the time domain, and therefore have negligible impact on perceived speech quality [ku2024explicit]. While some prior works [lee2023phaseaug, ku2024explicit] attempt to address these issues through data augmentation or specialized objectives, such methods often sacrifice compatibility with standard loss functions or require non-standard training pipelines.
@@
  • We propose EDNet, a versatile speech enhancement framework designed to handle diverse and compound distortions by embedding structural adaptability into its architecture.
@@ II-A Masking-based SE
Masking-based methods operate primarily in the time-frequency (TF) domain by learning masks that selectively suppress non-speech components while preserving speech-dominant regions. Early works like the Ideal Binary Mask (IBM) [hu2001speech, yilmaz2004blind, roman2011intelligibility, may2014generalization] apply hard decisions to each TF unit, significantly improving intelligibility but introducing artifacts due to coarse quantization [hummersone2014ideal, jin2009supervised]. To address this, Ideal Ratio Mask (IRM) [narayanan2013ideal, srinivasan2006binary, li2017ideal] was proposed, enabling soft masks with continuous values between 0 and 1 for smoother enhancement. The Spectral Magnitude Mask (SMM) [wang2014training] further relaxes the range constraint, allowing unbounded mask values and increasing modeling flexibility. As phase information gained importance, methods such as Phase-Sensitive Mask (PSM) [weninger2015speech, erdogan2015phase] and Complex Ideal Ratio Mask (cIRM) [williamson2015complex, kothapally2022skipconvgan, williamson2017time] were developed. Masking has been also extended to the latent feature space in time-domain architectures such as TasNet [luo2018tasnet], where masks are applied directly to learned feature representations, avoiding the constraints of the STFT domain. This feature-level masking has become central in many state-of-the-art models, particularly in the context of speech separation [subakan2021attention, zhao2023mossformer]. Despite their wide adoption, masking-based models remain fundamentally limited in scenarios involving severely missing spectral information, such as bandwidth limitation and packet loss [chao2025universal].
@@ II-B Mapping-based SE
Mapping-based speech enhancement methods aim to learn a nonlinear transformation that reconstructs clean speech directly from noisy input, typically in the waveform or spectral domain. Although masking remains the dominant approach in denoising, several studies have explored mapping-based alternatives. A representative example is SEGAN [pascual2017segan], which employs adversarial training for waveform mapping. Subsequent works increased generator capacity [phan2020improving] and combined adversarial loss with spectral-domain losses [pascual2019towards] for improved stability and fidelity. In dereverberation, mapping-based methods are motivated by the need of modeling long-range temporal smearing introduced by room reverberation. Early spectral regression models [han2015learning] used fully connected networks to predict clean spectra, followed by U-Net structures [ernst2018speech], and variants that replaced skip connections with multiple convolutional modules to improve feature propagation [kothapally2020skipconvnet]. More recently, deformable convolutional networks (DCN) [kothapally2024monaural] have been introduced context-sensitive transformations that extend beyond conventional masking and exhibit characteristics of both filtering and reconstruction-based approaches. Bandwidth extension is an area where mapping dominates, as it requires inferring missing high-frequency content. TF-domain methods [abel2018simple, li2015deep, lu2024towards] estimate high-band magnitude spectra from narrowband inputs and use heuristics or learned strategies for phase reconstruction. Time-domain approaches [kuleshov2017audio, wang2021towards] directly predict wideband waveforms, while vododer-based method such as NVSR [liu2022neural] uses a two-stage pipeline of spectral mapping followed by waveform generation via a neural vocoder. Despite their generative flexibility, mapping-based models can inadvertently distort clean input regions due to their global transformation behavior [nossier2020mapping, kumar2020nu, liu2022neural].
@@ II-C Hybrid Approachs
To overcome the individual limitations of masking and mapping approaches, several studies have explored hybrid architectures that combine the two, aiming to take advantage of their complementary strengths: masking offers selective noise suppression, while mapping enables detailed signal reconstruction. PhaseDCN [zhang2021deep] applies the hybrid approach with a focus on improving phase modeling to enhance denoising performance. It first generates a masked magnitude spectrum using an IRM and concatenates it with the noisy input. The combined features are fed into a complex-valued mapping network that produces a final output, implementing early stage fusion at the feature level. DBDIUNet [zhang2023dual] focuses on noisy and reverberant conditions by predicting a cIRM and a magnitude spectrogram through dual decoders, then averaging their outputs to produce the final result. The masking-and-inpainting [hao2020masking] method targets low signal-to-noise ratios (SNR) and non-stationary noise by first removing highly corrupted time-frequency regions using a binary mask, then reconstructing the removed region with a convolutional inpainting module. These two stages are executed sequentially, without an explicit fusion step. CMGAN [abdulatif2024cmgan], designed to handle various SE tasks such as denoising, dereverberation, and bandwidth extension, predicts magnitude masks and complex spectrograms in parallel and combines them using element-wise summation. Similarly, HD-DEMUCS [kim2023hd] targets general speech restoration using two parallel heterogeneous decoders, one for suppression and one for refinement, which are combined at the waveform level using a learnable weighting factor. This demonstrates the potential of hybrid methods in building general speech enhancement systems.
Despite differences in design and application, these approaches share several limitations. Most rely on fixed fusion strategies—such as averaging, concatenation, or summation—that are agnostic to input characteristics. While methods such as HD-DEMUCS introduce learnable fusion, the weighting mechanism operates only on the final decoder outputs, without direct guidance from the original input features. Moreover, these fusion mechanisms are typically applied globally at the output stage, lacking the ability to modulate the region-wise contributions of masking and mapping dynamically across time and frequency. These observations underscore the need for more flexible and sophisticated integration strategies that can coordinate masking and mapping in a input content-aware, task-agnostic manner.
@@ II-D Phase Modeling
Early speech enhancement models primarily focused on magnitude estimation [xu2014regression, ai2019dnn], often reusing the noisy phase during synthesis. However, it is now widely recognized that accurate phase reconstruction is essential for achieving high perceptual quality [paliwal2011importance]. One line of research addresses phase prediction by directly operating in the complex domain either through masking [hao2021fullsubnet, choi2018phase, zhao2022frcrn] or mapping [wang2023tf, wang2020complex], where both magnitude and phase are implicitly modeled via real and imaginary components. While conceptually straightforward, this approach often suffers from optimization instability and limited interpretability, especially when paired with magnitude-focused objectives such as spectral magnitude loss or mask estimation targets [lu2023explicit]. To improve the stability and effectiveness of phase modeling, recent studies [chao2024investigation, lu2023explicit, lu2024towards] have moved toward explicitly decoupling magnitude and phase and modeling them through separate branches.
Supervision for phase modeling varies by modeling approach. In complex-valued formulations, losses are typically applied in terms of complex spectral distance. In contrast, explicit phase prediction often requires trigonometric losses [takamichi2018phase] or anti-wrapping functions [ai2023neural] to handle phase discontinuities, and is frequently combined with derivative-based objectives such as group delay (GD) or instantaneous angular frequency (IAF) [zhang2024unrestricted, masuyama2020phase]. Most of these methods, however, assume a single ground-truth (GT) phase, implicitly enforcing strict alignment, which can penalize perceptually irrelevant shifts and destabilize learning [ku2024explicit].
While the limitations of strict phase alignment have begun to attract attention, only a few studies have explored ways to relax this constraint. PhaseAug [lee2023phaseaug], originally proposed for vocoder training, introduces random phase shifts during training to expose the model to multiple plausible alignments. However, the loss is still computed with respect to a single shifted target, meaning the model is implicitly encouraged to align with the shifted version of the GT. Ku et al. [ku2024explicit] take a different approach by removing phase GT entirely and using magnitude-phase consistency as an indirect supervision signal. While this avoids explicit alignment, it prevents the use of standard loss functions—such as waveform or complex spectral losses—since applying them would implicitly reintroduce supervision toward a fixed target phase. These issues suggest the need for a supervision strategy that resolves alignment ambiguity while preserving compatibility with existing loss functions and training pipelines.
@@ III-A Dual-stream Architecture
EDNet is designed with a dual-stream architecture to separately process magnitude and phase spectrograms, as shown in  Fig. 1. Instead of using a shared encoder, it employs separate dilated DenseNet encoders [lu2023explicit] with feature channel dimensions of 64 and 32 for magnitude and phase, respectively. This separation is intended to allow magnitude and phase, which have distinct characteristics—magnitude follows a linear scale, where phase exhibits periodicity—to be processed individually, aiming to prevent interference between their estimations. The decoder also follows this structure, using separate dilated DenseNet decoders [lu2023explicit] for magnitude and phase. The magnitude decoder directly reconstructs the spectrogram using ReLU activation [glorot2011deep], as feature-wise masking is handled within the GM module.
The input speech signal is first transformed using Short-Time Fourier Transform (STFT), decomposing it into magnitude and phase spectrograms. The magnitude spectrogram undergoes power-law compression to enhance feature learning. Each spectrogram is then encoded separately using the respective encoders. Noisy magnitude features are progressively refined through three stacked GM modules. The enhanced magnitude features are then concatenated with the noisy phase features, reduced to 32 channels via a 1×1 convolution, and processed through three stacked Time-Frequency (TF) Mamba blocks [chao2024investigation], which model long-range dependencies across temporal and spectral domains using bidirectional Mamba [gu2024mambalineartimesequencemodeling], to refine the phase. Finally, the enhanced spectrograms are reconstructed using the decoders, and the enhanced waveform is obtained by combining them with Inverse STFT (ISTFT).
@@ III-C1 Phase Shift Estimation
where $f_{AW}$ is an anti-wrapping function that ensures phase differences remain within a range $\pi,\pi]$. However, the presence of the anti-wrapping function introduces non-linearity, making it difficult to derive a closed-form solution for the optimal $n$.
@@
This formulation allows us to approximate the optimal shift by assuming a locally linear relationship between the anti-wrapped phase difference and the shift term $2\pi fn/N$. For a restricted range of $n$ in [0.5,0.5][-0.5,0.5], where shift term remains within the anti-wrapping range, this approximation holds effectively, enabling a differentiable solution. Taking the derivative with respect to $n$ and solving for zero, we obtain: